# Robust Containment Queries over Collections of Rational Parametric Curves via Generalized Winding Numbers

### v1

@@ 4. Generalized Winding Numbers for Curved Geometry
Most importantly, the winding number of such a closure can always be computed exactly without the need to appeal to quadrature by Equation 4, as the angle subtended by the introduced linear segment can be evaluated through a (relatively) inexpensive arccosine. In this way, we can always solve the problem of the generalized winding number using a solution to the integer winding number problem on the closed curve, as subtracting away the contribution of this linear closure can be done independently of the original curve geometry. While we present our own algorithm to compute this integer winding number in Section 5 that outperforms known alternatives, we note that this strategy is fully compatible with more conventional techniques for computing containment queries in closed regions bounded by curves, e.g. Bézier clipping (Nishita et al., 1990).
@@
This principle is particularly useful in the case when $w_{\Gamma\cup C}(q)=0$, i.e. the query point is located outside the closed shape $\Gamma\cup C$, or outside the convex hull of $\Gamma$ itself, where we have $w_{\Gamma}(q)=-w_{C}(q)$. This usage is introduced in Jacobson et al. (2013), where it is used in a hierarchical evaluation of the winding number for collections of linear triangular facets. An important consequence of this for curved geometry is that for query points that are far enough away, we can reverse the orientation of $C$ and treat the shape as equivalent to its straight line closure, and compute the winding number of the entire curved segment immediately and exactly.
We further extend this principle to evaluate generalized winding numbers of points that are arbitrarily close to the curve with the same assurances of exact accuracy. In such cases, $\Gamma$ can be replaced with a piecewise linear approximation $\widetilde{\Gamma}$ so long as the integer winding number at $q$ remains unchanged between $\Gamma\cup C$ and $\widetilde{\Gamma}\cup C$ (see Figure 9). Doing so necessarily leaves the generalized winding number at $q$ unchanged as well, as we have
@@ 5. Winding Number Algorithms
  • Otherwise, we bisect the Bézier curve and repeat the algorithm on each half. As a base case, we check whether the curve is approximately linear (Algorithm 3), as such segments can be added directly to the polyline.
@@
In practice, because Bézier curves are smooth nearly everywhere on their interior, the recursive bisection step needs to be done relatively few times before the query point is found to be outside the convex hulls of each component, even for points very close to the curve. However, it can be problematic if the query point is located directly on the curve, as in such cases the winding number is not formally defined. Since we need our algorithm to be robust to such input points, we discuss this case extensively in Section 6.
@@
Despite the simplicity of the above procedure, directly computing containment within a Bézier curve’s convex hull can be prohibitively expensive. Instead, we first determine if the query is contained in an axis-aligned bounding box that encompasses the convex hull. Although this can have a much larger area, the bounding box containment query is inexpensive and most points will be exterior to it. If it is not, we use a test from Ying and Hewitt (2003) to check if the control polygon (defined by a Bézier curve’s control nodes) is already simple and convex (see Algorithm 4). If so, we perform the point-in-polygon test to easily determine if the query lies outside the convex hull of the control polygon. For efficiency, we note that when a curve is simple and convex, its subcomponent curves will be as well.
Input: $\Gamma$: Rational Bézier curve $\Gamma$
@@
$C\leftarrow(1-t)\Gamma(1)+t\Gamma(0)$
@@
ALGORITHM 1 WindingNumberCurve: Evaluate the generalized winding number for an arbitrary rational Bézier curve
Input: $\Gamma\cup C$: closed rational Bézier curve
@@
       $C_{0}\leftarrow(1-t){\Gamma_{0}(0)}+t{\Gamma_{0}(1)}$
@@
ALGORITHM 2 IntegerWindingNumberCurve Evaluate the integer winding number for a rational Bézier curve closed by a linear segment. We use an algorithm from Hormann and Agathos (2001) for PolygonWindingNumber.
Input: $\Gamma$: Rational Bézier curve
@@
$C_{0}\leftarrow(1-t){\Gamma_{0}(0)}+t{\Gamma_{1}(1)}$
@@
ALGORITHM 3 isApproximatelyLinear: Return true if each control point of a rational Bézier curve is within a given tolerance of the closure.
Input: $\Gamma$: Rational Bézier curve with control nodes ${P_{0}},\cdots,{P_{n}}$
@@
ALGORITHM 4 isSimpleConvex: Return true if the polygon of control points of a rational Bézier curve is simple and convex (Adapted from Ying and Hewitt (2003))
@@ 6. Generalized Winding Numbers for Coincident Points
The generalized winding number is, in a strict mathematical sense, undefined for points located directly on the curve. This poses a practical problem during the implementation of a containment query, as they are often executed massively in parallel for large clusters of points without any a priori knowledge of their position relative to the boundary. Furthermore, our desire for exact geometric fidelity precludes us from applying the standard trick of perturbing such points randomly to place them definitively on one side of the singular boundary. In spite of this, it is necessary for robustness with respect to downstream applications that the algorithm definitively return a value that adheres to a convention that is mathematically justified and intuitive to the caller of WindingNumberCurve.
@@
Despite this conceptual clarity, evaluating coincident winding numbers is complicated in practice, as it is computationally expensive to perform the projection necessary to identify when a point is located exactly on a Bézier curve (Ying and Hewitt, 2003). The exception to this is the endpoints, which are interpolated exactly by the first and last control points of the curve. At such points, the coincident winding number is equal to the signed angle spanned by the non-coincident endpoint, and a tangent vector at the coincident endpoint (see Figure 11). Furthermore, computing such a tangent is trivial at the endpoint of a Bézier curve, as it is defined by the endpoint and the adjacent control node. Interior points become endpoints in linear time by the repeated bisections of our algorithm. As this spanned angle must account for full revolutions of this span around the tangent vector, our algorithm only applies this edge-case to curves with simple and convex control polygons (see Algorithm 5). This particular criterion is typically met after very few bisections, and ensures that no full revolutions can occur on the local subcurve.
@@
Input: $\Gamma$: Rational Bézier curve
@@ 7.1. Robustness of Containment Queries on Curved Geometry
We first show the utility of a generalized winding number approach to containment queries in the context of messy CAD geometry. We compare our approach to conventional ray casting for containment, which remains the de-facto standard for exact containment in curved watertight geometry. Consider the geometry of Figure 12, which is composed of 87 linear segments and 477 cubic Bézier curves. We remove the marked curve in Figure 12(left) and determine containment at each pixel in the image using a simple ray casting algorithm that extends the ray to the right of the pixel and counts the number of intersections with the shape. As expected, the use of a ray casting algorithm causes catastrophic issues, in the sense that the containment errors are located at a great distance from the actual geometric error. This makes the model unusable in the applications of interest without specific oversight and (often user-driven) correction of these errors.
@@
We then apply our generalized winding number algorithm to the same geometry in Figure 13, evaluating at each pixel the winding number field generated by the bounding geometry. We see that the deletion of the same curve does impact the field, but that this impact is localized to the site of the geometric error. In some sense, the locality of this degradation is unintuitive, as containment is necessarily a global property of points relative to bounding geometry, i.e. the winding number of a point is dependent on every curve. Importantly, however, this influence of distant curves becomes rapidly negligible. While the exact value is now a non-integer value (nearly) everywhere in the domain, away from the deleted curve the difference from the nearest integer is nowhere large enough to have an adverse effect on the actual determination of containment. Thus, points which we expect to be interior have winding numbers close to 1, while those of points we expect to be exterior are close to 0.
In these examples, we decide containment from the scalar field of fractional values by rounding each to the nearest integer and applying a non-zero rule. In Figure 13(right) we see that the rounded winding number produces a clear boundary between regions along the 1/2 isocurve. While this implied boundary does not have the same curved geometry of the deleted shape, it is a reasonable approximation and nevertheless allows for the surface to be used immediately in applications. This behavior is especially desirable when geometric errors arise from small gaps between components of the mesh, as the resulting containment queries reflect the simplest closure of the shape permitted by the provided geometric data.
Furthermore, this is not the only way one can determine containment from a fractional winding number. Access to the full winding number field through our algorithm ultimately results in a standard segmentation problem, which can be solved with any number of strategies. For example, Jacobson et al. (2013) meshes the interior of the shape, utilizing an energy minimization technique to enforce additional smoothness in the resulting segmentation. Because we are primarily concerned with efficient evaluation of the winding number more generally, we do not explore additional segmentation techniques in this work.
@@
As noted in Jacobson et al. (2013), a useful perspective on the fractional value of the winding number is as a measure of confidence in the derived containment query. This “confidence” is not meant to be interpreted in a rigorous statistical sense, as the true intent of a mesh can never be known in principle. Nevertheless, a winding number of 0.99 indicates that the query point is “more likely” to have been intended as an interior point than a winding number of 0.75, even if both points are classified identically. In this way, points on the winding number field with values that are close to half-integers tend to correlate with the location of geometric inaccuracy. Through this extension to curved geometry, we can now extend this confidence framework to a much broader class of geometric models.
@@
To explore this further, we consider Figure 14, which features a shape that is intentionally deformed such that no pairs of adjacent edges are connected. From a distance. i.e. the scale at which the model designed could have determined the geometry to be “good enough,” there are functionally no irregularities in the shape’s appearance. This mirrors the realistic setting, in which such imperceptible tolerances remain unexpected and catastrophic for conventional containment algorithms. In contrast, the generalized winding number is indifferent to these errors – even at this frequency the degradation of the field remains localized to the meshing error. As before, we apply our algorithm to each pixel in the bounding region, and see that despite the existence of geometric and topological errors throughout the image, the resulting values still very clearly partition the shape into the expected interior and exterior, even at nearer scales.
To explore this further, we identify in this image each “uncertain” classification, which we heuristically consider to be each point with a winding number in the range [0.25, 0.75]. Otherwise, we can be reasonably confident in the determination of containment. Plotting the difference of the generalized winding number field from the nearest integer reveals a number of its important characteristics, and their impact on downstream applications that depend on containment queries. First, even small geometric errors have far-reaching influence on the scalar field itself, as seen by how few points in the image have a winding number that is within two decimal places of an integer. This means that while the winding number field degrades gracefully away from geometric errors, very rarely does it decay completely to zero. Close to these errors, however, the field decays rapidly, and only in a very small region is the winding number different enough from an integer so as to indicate doubt in the classification. Thus, for the vast majority of points, the classification will be done with near certainty.
@@ 7.2. Algorithm Performance
We now consider the computational performance of Algorithms 1 and 2 through comparison to state-of-the-art techniques adapted to collections of parametric curves.
@@
Perhaps the most common winding number-based approach for watertight geometry would be to discretize the shape into approximating polygons with linear edges and to apply a standard point-in-polygon algorithm to the polygonal shapes to determine containment. If watertightness cannot be guaranteed, the generalized winding number approaches of Jacobson et al. (2013); Trettner et al. (2022) admit a straightforward adaptation to a linearized shape.
@@
In demonstration of this fact, consider Figure 15. To begin, we note that the shape itself is non-watertight and non-manifold, making it a particularly challenging case for conventional containment queries. Nevertheless, the fractional winding number field clearly partitions the space according to an intuitive understanding of its interior. We compare refinements of the curves in the shape into a fixed number of linear segments against an approximate likelihood that points are ultimately misclassified. We can see that although the likelihood of misclassification decreases with increased linear refinement, it does so at a rate that we consider to be far too slow to be practical for many applications of interest, especially for downstream applications that are sensitive to any misclassifications. This is also an issue for methods that sample the bounding geometry non-deterministically, such as those of Nooruddin and Turk (2003); Barill et al. (2018); Sawhney and Crane (2020). In their place, we turn our attention to methods that operate directly on curved geometry without approximation.
@@
Figure 15. Given a non-watertight, non-manifold shape, we observe the practical effect of curve linearization on geometric accuracy. We consider various linearizations of the curve shapes, and record how many points out of $10^{5}$ randomly sampled from a uniform distribution in the domain are misclassified. We note that even at high levels of refinement there are still a considerable number of misclassifications, which is unacceptable in many downstream applications of interest. In contrast, the proposed method is able to achieve perfect geometric fidelity for points that are arbitrarily close to the boundary.
@@
There remains the issue that the methods known to the authors that operate directly on curved geometry à la Bézier clipping (Nishita et al., 1990) can do so only in the watertight case. This regrettably means that no direct comparison for the performance Algorithm 1 can be made at this time. In its place, we note that our procedure for computing the generalized number for any open, curved object requires applying an arbitrary integer winding number algorithm on the closed shape and subsequently subtracting out the contribution of the closure of the curve. It is in this context that we demonstrate the performance of Algorithm 2, which is one such algorithm that computes the integer winding number of a closed, curved object. We show that Algorithm 2 is considerably more performant than alternatives utilizing crossing numbers for containment queries on closed, curved geometry, as such methods necessarily introduce inefficiencies through a reliance on ray casting.
@@
We first consider a common procedure for ray casting that applies a geometric binary search to compute ray-Bézier curve intersections (Farin, 2001). In this algorithm, we extend from the query a ray in the direction of the closest edge of a bounding box, and recursively bisect the curve until guarantees for potential intersections can be made. If a subcurve is approximately linear (see Algorithm 3), then the signed intersection can be recorded. If the ray does not intersect the subcurve’s bounding box, then it can be discarded. This procedure converges to points of intersection linearly, and is capable of identifying multiple intersections with the same curve.
@@
To make the appropriate comparison to Algorithm 2 in the context of nonwatertight geometry, we utilize each within the framework described in Algorithm 1. That is to say, we iterate over each open curve in the model and close it, compute the integer winding number with Algorithm 2 or one of the above alternate techniques, and subtract out the contribution of the closure. We explicitly note that, while these alternate techniques themselves are well-known in the context of ray-tracing and containment of closed shapes, computing generalized winding numbers this way is, to our knowledge, in itself a novel application of their use.
@@
Our first example in Figure 16 uses a relatively simple shape featuring at most polynomial, cubic Bézier curves. To compare the three algorithms (binary search, Bézier clipping, and the proposed approach presented in Algorithm 2), we sample 250,000 points from a bounding box for the shape using a uniform distribution, and count the number of curve evaluations on each that needs to be made to determine containment with perfect geometric fidelity. Again, we emphasize that all three methods are being used here to compute the generalized winding number, which they do by computing the integer winding number of the closed shape for each curve in the model.
@@
Figure 16. We compare the cost of a conventional geometric bisection method, Bézier clipping, and our proposed method on a watertight model. We uniformly sample 250,000 points from a bounding box and measure the number CAD model evaluations, i.e. curve subdivisions, that are necessary to evaluate the winding number at each point to full geometric precision. In the histogram, we omit the number of sampled points that require zero function evaluations, as this occurs whenever the point is outside each curve’s bounding box, occurring at the same frequency for each method.

### v2

@@ 4. Generalized Winding Numbers for Curved Geometry
Most importantly, the winding number of such a closure can always be computed exactly without the need to appeal to quadrature by Equation 4, as the angle subtended by the introduced linear segment can be evaluated through a (relatively) inexpensive arccosine. In this way, we can always solve the problem of the generalized winding number using a solution to the integer winding number problem on the closed curve, as subtracting away the contribution of this linear closure can be done independently of the original curve geometry. While we present our own algorithm to compute this integer winding number in Section 5 that outperforms known alternatives, we note that this strategy is fully compatible with more conventional techniques for computing containment queries in closed regions bounded by curves, e.g., Bézier clipping (Nishita et al., 1990).
@@
[…]
Figure 7: A visual demonstration of the generalized winding number is shown for a simple curve. The winding number field for the open curve is unknown, but the field for the closed shape is integer valued, and the field for the closure is known exactly.
This principle is particularly useful in the case when $w_{\Gamma\cup C}(q)=0$, i.e., the query point is located outside the closed shape $\Gamma\cup C$, or outside the convex hull of $\Gamma$ itself, where we have $w_{\Gamma}(q)=-w_{C}(q)$. This usage is introduced in Jacobson et al. (2013), where it is used in a hierarchical evaluation of the winding number for collections of linear triangular facets. An important consequence of this for curved geometry is that for query points that are far enough away, we can reverse the orientation of $C$ and treat the shape as equivalent to its straight line closure, and compute the winding number of the entire curved segment immediately and exactly.
We further extend this principle to evaluate generalized winding numbers for points that are arbitrarily close to the curve with the same assurances of exact accuracy. In such cases, $\Gamma$ can be replaced with a piecewise linear approximation $\widetilde{\Gamma}$ so long as the integer winding number at $q$ remains unchanged between $\Gamma\cup C$ and $\widetilde{\Gamma}\cup C$ (see Figure 8). Doing so necessarily leaves the generalized winding number at $q$ unchanged as well, as we have
@@
[…]
Figure 8: A visual demonstration of the principle that the winding number field generated by a curve is closely related to that of its linear closure, shown on four increasingly simple curves which share endpoints.
@@ 5. Winding Number Algorithms
  • Otherwise, we bisect the Bézier curve and repeat the algorithm on each half. As a base case, we check whether the curve is approximately linear (Algorithm 3) as such segments can be added directly to the polyline.
@@
To improve computational efficiency in the case where subdivision is expensive and exact geometric accuracy is not needed, we optionally provide in Algorithm 3 an early stopping criterion. We choose this particular heuristic because it is easier to evaluate than the distance between the query point and the curve, and has a more intuitive spatial interpretation than a simple bound on the maximum number of curve subdivisions. In practice, Bézier curves are smooth nearly everywhere on their interior, and so the recursive bisection step needs to be done relatively few times before the query point is found to be outside the convex hulls of each component, even for points very close to the curve. Importantly, this causes the accuracy of Algorithm 1 and runtime of the procedure to both be reasonably insensitive to this numerical tolerance, as the algorithm is likely to terminate for these smooth curves well before there has been sufficient refinement for subdivisions to be considered linear. Indeed, in the numerical examples we provide, we set this value to machine epsilon and find no significant loss in performance (See Section 7.2).
We note that it can be problematic if the query point is located directly on the curve, as in such cases the winding number is not formally defined. Since we need our algorithm to be compatible to such input points, we discuss this case extensively in Section 6.
@@
[…]
Figure 9: Fully described in the caption.
Despite the simplicity of the above procedure, directly computing containment within a Bézier curve’s convex hull can be prohibitively expensive. Instead, we first determine if the query is contained in an axis-aligned bounding box that encompasses the convex hull. Although this can have a much larger area, the bounding box containment query is inexpensive and most points will be exterior to it. If it is not, we use a test from Ma and Hewitt (2003) to check if the control polygon (defined by a Bézier curve’s control nodes) is already simple and convex (see Algorithm 4). If so, we perform the point-in-polygon test to determine if the query lies outside the convex hull of the control polygon. For efficiency, our implementation notes that when a curve is simple and convex, its subcomponent curves will be as well.
Input: $\Gamma$: Rational parametric Bézier curve $\Gamma$
@@
$C(t)\leftarrow(1-t)\Gamma(1)+t\Gamma(0)$
@@
ALGORITHM 1 WindingNumberCurve: Evaluate the generalized winding number for an arbitrary rational parametric Bézier curve
Input: $\Gamma\cup C$: Closed rational parametric Bézier curve
@@
       $C_{0}(t)\leftarrow(1-t){\Gamma_{0}(0)}+t{\Gamma_{0}(1)}$
@@
ALGORITHM 2 IntegerWindingNumberCurve Evaluate the integer winding number for a rational parametric Bézier curve closed by a linear segment. We use an algorithm from (Hormann and Agathos, 2001) for PolygonWindingNumber.
Input: $\Gamma$: Rational parametric Bézier curve
@@
$C_{0}(t)\leftarrow(1-t){\Gamma_{0}(0)}+t{\Gamma_{1}(1)}$
@@
ALGORITHM 3 isApproximatelyLinear: Return true if each control point of a rational parametric Bézier curve is within a given tolerance of the closure.
Input: $\Gamma$: Rational parametric Bézier curve with control nodes ${P_{0}},\cdots,{P_{n}}$
@@
ALGORITHM 4 isSimpleConvex: Return true if the polygon of control points of a rational parametric Bézier curve is simple and convex (Adapted from (Ma and Hewitt, 2003))
@@ 6. Generalized Winding Numbers for Coincident Points
The generalized winding number is, in a strict mathematical sense, undefined for points located directly on the curve. This poses a practical problem during the implementation of a containment query, as they are often executed massively in parallel for large clusters of points without any a priori knowledge of their position relative to the boundary. Furthermore, our desire for exact geometric fidelity precludes us from applying the standard trick of perturbing such points randomly to place them definitively on one side of the singular boundary. In spite of this, it is necessary for compatibility with respect to downstream applications that the algorithm definitively return a value that adheres to a convention that is mathematically justified and intuitive to the caller of WindingNumberCurve.
@@
Despite this conceptual clarity, evaluating coincident winding numbers is complicated in practice, as it is computationally expensive to perform the projection necessary to identify when a point is located exactly on a Bézier curve (Ma and Hewitt, 2003). The exception to this is the endpoints, which are interpolated exactly by the first and last control points of the curve. At such points, the coincident winding number is equal to the signed angle spanned by the non-coincident endpoint, and a tangent vector at the coincident endpoint (see Figure 10). Furthermore, computing such a tangent is trivial at the endpoint of a Bézier curve, as it is defined by the endpoint and the adjacent control node. Interior points become endpoints in linear time by the repeated bisections of our algorithm. As this spanned angle must account for full revolutions of this span around the tangent vector, our algorithm only applies this edge-case to curves with simple and convex control polygons (see Algorithm 5). This particular criterion is typically met after very few bisections, and ensures that no full revolutions can occur on the local subcurve.
@@
[…]
Figure 10: Fully described by the caption.
Input: $\Gamma$: Rational parametric Bézier curve
@@ 7.1. Robustness of Containment Queries on Curved Geometry
We first show the utility of a generalized winding number approach to containment queries in the context of messy CAD geometry. We compare our approach to conventional ray casting for containment, which remains the de-facto standard for exact containment in curved watertight geometry. Consider the geometry of Figure 11, which is composed of 87 linear segments and 477 cubic Bézier curves. We remove the marked curve in Figure 11 and determine containment at each pixel in the image using a simple ray casting algorithm that extends the ray to the right of the pixel and counts the number of intersections with the shape. As expected, the use of a ray casting algorithm causes catastrophic issues, in the sense that the containment errors are located at a great distance from the actual geometric error. This makes the model unusable in the applications of interest without specific oversight and (often user-driven) correction of these errors.
@@
[…]
Figure 11: A butterfly vector graphic made up of 500+ curves is shown. When one curve is deleted, the ray casting algorithm produces far-reaching errors in containment queries.
We then apply our generalized winding number algorithm to the same geometry in Figure 12, evaluating at each pixel the winding number field generated by the bounding geometry. In this example, points which we expect to be interior to this shape have winding numbers close to 1, while the value for points we expect to be exterior is close to 0. We see that deleting the curve does degrade the field, but that the most severe effects are localized to the site of the geometric error. In some sense, the locality of this degradation is unintuitive, as containment is necessarily a global property of points relative to bounding geometry, i.e., the winding number at every point is dependent on every curve. Importantly, however, the influence of distant curves becomes rapidly negligible. While the exact value of the winding number now a non-integer value (nearly) everywhere in the domain, away from the deleted curve the difference from the nearest integer is nowhere large enough to have an adverse effect on the actual determination of containment.
A natural mapping of winding numbers to containment classifications is to round each value to the nearest integer and apply a non-zero rule. For example, in Figure 12 we see that the rounded winding number produces a clear boundary between regions along the 1/2 isocurve. While this implied boundary lacks the curvature of the deleted curve, it is a reasonable approximation and nevertheless allows for the surface to be used immediately in applications. This behavior is especially desirable when geometric errors arise from small gaps between components of the mesh, as the resulting containment query reflects the simplest closure of the shape permitted by the provided geometric data.
@@
[…]
Figure 12: The same butterfly graphic from Figure 11 is shown, but the generalized winding number field is used to determine containment instead. The field degrades smoothly around the deleted curve, and rounding the winding number produces an approximated boundary between interior and exterior regions.
@@
Through Figure 13, we reiterate the two important kinds of robustness in this work. The shape in this example is intentionally deformed such that no pairs of adjacent edges are connected. From a distance, i.e., the scale at which the model geometry is visually “good enough,” there are functionally no irregularities in the shape’s appearance. This mirrors the realistic setting, in which such imperceptible tolerances are unexpected and remain catastrophic for conventional containment algorithms. As before we apply our algorithm to each pixel in the bounding region, and plot the winding number field alongside its difference from the nearest integer.
The first type of robustness is derived from our specific computation of the generalized winding number through Algorithm 1, which is completely robust to non-watertight geometry. Because we only ever consider individual curves without any reference to their overall arrangement, both the computational performance and accuracy of our calculation is completely unaffected by the non-watertightness of the shape. Furthermore, our adaptive algorithm ensures that the winding number is computed stably at arbitrary distances from individual curves. On the other hand, any containment query derived from generalized winding numbers is robust to geometric artifacts in the boundary. This is because although even small errors have far-reaching influence on the scalar field of winding numbers, this influence degrades rapidly and gracefully away from the sources of these errors. Even at this frequency of topological errors in this example, the resulting values of the field still very clearly partition the shape into the expected interior and exterior.
It is noted in Jacobson et al. (2013) that a useful perspective on the fractional value of the winding number is as a measure of confidence for the derived containment query. For example, points for which the winding number is close to an integer can be considered “more likely” to be classified accurately according to the unknown, but presumably watertight shape that lacks any of the present boundary artifacts. On the other hand, points with winding numbers near half-integer values are often close to the boundaries that are only implied by the rounding heuristic, and correlate with containment classifications that are less informed by existing surface elements. In Figure 13, we specifically identify each point in the image with a winding number in the range [0.25, 0.75], which are the only points that we can therefore consider to have an “uncertain” classification according to our rounding heuristic. From this, we can see that the vast majority of points in this image are classified with high confidence.
Of course, the rounding heuristic is not the only way one can determine containment from a fractional winding number, as access to the full winding number field through our algorithm permits treatment of containment through standard segmentation strategies. For example, Jacobson et al. (2013) meshes the interior of the shape, utilizing an energy minimization technique to enforce additional smoothness in the resulting segmentation. Additionally, the above usage of the winding numbers as a measure of confidence can be made more rigorous through further statistical analysis. In Sellán and Jacobson (2022), the generalized winding number field is used as a prior for a number of statistical distributions, including likelihood of containment. This rigor is desirable for certain applications, as the fractional value of the winding number alone cannot be interpreted as a probability for the accuracy of the derived containment query. Because we are ultimately concerned with efficient evaluation of the winding number in this work, we use only the rounding heuristic to determine containment in our examples.
@@
[…]
Figure 13: A very-slightly non-watertight fish shape is shown with the generalized winding number field computed for each pixel. The absolute difference in log scale between the computed winding number and the rounded winding number is shown, with points that have a difference greater than 0.25 highlighted. These points occur only in small radii around gaps in the shape.
@@ 7.2. Algorithm Performance
We now consider the computational performance of Algorithms 1 and 2. In each of the following examples, we evaluate our algorithm in the context where exact geometric fidelity is prioritized, and set all numerical tolerances to zero. For comparison, we consider state-of-the-art techniques that have been adapted to collections of parametric curves. Perhaps the most common winding number-based approach for watertight geometry would be to discretize the shape into approximating polygons with linear edges and to apply a standard point-in-polygon algorithm to the polygonal shapes to determine containment. If watertightness cannot be guaranteed, the generalized winding number approaches of Jacobson et al. (2013); Trettner et al. (2022) admit a straightforward adaptation to a linearized shape.
In demonstration of this fact, consider Figure 14, where we compare the performance of our serial implementation of the proposed adaptive strategy to such a linearized method. Specifically, we consider refinements of each curves in the shape into a fixed number of linear segments. We note that the shape considered is non-watertight and non-manifold, making it a particularly challenging case for conventional containment queries. Nevertheless, the fractional winding number field clearly partitions the space according to an intuitive understanding of its interior.
By comparing wall-clock runtime, we see that the use of a fixed level of refinement of the curve leads to a more efficient calculation of the generalized winding number as expected, as curve subdivisions can be computed during a preprocessing step and reused. While the analogous overhead in the adaptive algorithm can be similarly mitigated through hash maps that cache the results of curve subdivisions, we consider further implementation optimizations for collections of curves to be future work, and instead focus here on efficient and accurate calculation for individual curves. Indeed, when geometric fidelity is not a concern, using a fixed linearization to compute generalized winding numbers with Algorithm 1 is a reasonable approach, although it may involve a considerable memory footprint for larger models.
@@
Importantly, such a linearized method offers no guarantees for geometric fidelity. By comparing the level of refinement against an approximate likelihood that points are ultimately misclassified, we see that although the rate of misclassification decreases with increased linear refinement, it does so at a rate that we consider to be far too slow to be practical for many applications of interest, especially for downstream applications that are sensitive to any misclassifications. This is also an issue for methods that sample the bounding geometry non-deterministically, such as those of Nooruddin and Turk (2003); Barill et al. (2018); Sawhney and Crane (2020). In contrast, the proposed method is able to achieve perfect geometric fidelity for points that are arbitrarily close to the boundary. In place of these fixed linearization methods, we turn our attention to those that operate directly on curved geometry without approximation.
@@
Figure 14. Given a non-watertight, non-manifold shape, we observe the practical effect of curve linearization on geometric accuracy and runtime. We compare our adaptive strategy to those that uses a uniform, piecewise linearization of the curve at fixed levels of refinement. We record the total time spent to evaluate the generalized winding number for $10^{5}$ points randomly sampled from a uniform distribution, and the number of these points which are ultimately misclassified. We see that a fixed linearization can offer better computational performance, as subdivisions of the curve can be efficiently precomputed. However, we note that even at high levels of refinement there are still a considerable number of misclassifications, which is unacceptable in many downstream applications of interest.
[…]
Figure 14: A per-point comparison of the performance of the adaptive strategy to a fixed linearization method is shown. The total wall-clock time needed to evaluate the generalized winding number for all points using various levels of a fixed linearization is shown, along with the number of misclassifications induced by each. The adaptive strategy is slower, particularly for points that are near the curve. However, every level of linearization used introduces a non-zero number of misclassifications.
@@
There remains the issue that the methods known to the authors that operate directly on curved geometry à la Bézier clipping (Nishita et al., 1990) can do so only in the watertight case. This regrettably means that no direct comparison for the performance Algorithm 1 can be made at this time. In its place, we note that our procedure for computing the exact generalized winding number for any open, curved object requires applying an arbitrary integer winding number algorithm on the closed shape and subsequently subtracting out the contribution of the closure of the curve. It is in this context that we demonstrate the performance of Algorithm 2, which is one such algorithm that computes the integer winding number of a closed, curved object. We show that Algorithm 2 is considerably more performant than alternatives utilizing crossing numbers for containment queries on closed, curved geometry, as such methods necessarily introduce inefficiencies through a reliance on ray casting.
@@
We first consider a common procedure for ray casting that applies a geometric binary search to compute ray-Bézier curve intersections (Farin, 2001). In this algorithm, we extend from the query a ray in the direction of the closest edge of a bounding box, and recursively bisect the curve until guarantees for potential intersections can be made. If a subcurve is sufficiently linear (see Algorithm 3) such that an intersection is guaranteed, then the signed intersection can be recorded. If the ray does not intersect the subcurve’s bounding box, then it can be discarded. This procedure converges to points of intersection linearly, and is capable of identifying multiple intersections with the same curve.
@@
To make the appropriate comparison to Algorithm 2 in the context of exact winding numbers for non-watertight geometry, we utilize each within the framework described in Algorithm 1. That is to say, we iterate over each open curve in the model and close it, compute the integer winding number with Algorithm 2 or one of the above alternate techniques, and subtract out the contribution of the closure. We note that, while these alternate techniques themselves are well-known in the context of ray-tracing and containment of closed shapes, computing generalized winding numbers in this way is, to our knowledge, itself a novel application of their use.
@@
Our first example in Figure 15 uses a relatively simple shape featuring at most polynomial, cubic Bézier curves. To compare the three algorithms (binary search, Bézier clipping, and the proposed approach presented in Algorithm 2), we randomly sample 250,000 points from a bounding box around the shape using a uniform distribution, and count the number of curve evaluations on each that needs to be made to determine containment with perfect geometric fidelity. Again, we emphasize that all three methods are being used here to compute the generalized winding number, which they do by computing the integer winding number of the closed shape for each curve in the model.
@@
Figure 15. We compare the cost of a conventional geometric bisection method, Bézier clipping, and our proposed method on a watertight model. We uniformly sample 250,000 points from a bounding box and measure the number CAD model evaluations, i.e., curve subdivisions, that are necessary to evaluate the winding number at each point to full geometric precision. In the histogram, we omit the number of sampled points that require zero function evaluations, as this occurs whenever the point is outside each curve’s bounding box, occurring at the same frequency for each method.
[…]
Figure 15: Given a watertight rose shape, a heatmap of the number of curve evaluations needed to evaluate the generalized winding number for 250,000 points in the domain is shown for each of a binary search method, Bezier clipping, and the proposed method. The proposed method requires the fewest curve evaluations for the vast majority of points. A histogram shows the distribution of curve evaluations needed for each method.